its a right triangle side A=X, side B=12 and the hypotenuse=6+X and im supposed to solve for x but when i set it up the x's keep cancelling out and equaling zero, how do i solve this, please help
Answer by mathhelp@(10)

its a right triangle side A=X, side B=12 and the hypotenuse=6+X and im supposed to solve for x but when i set it up the x's keep cancelling out and equaling zero, how do i solve this, please help
use the formula:
a^2+b^2=c^2
so your a = x
b=12
c=6+x
x^2+144=(6+x)^2
expand:
x^2+144=x^2+12x+36
x^2+108=x^2+12x
cancel out the x^2
108=12x
x=9
so your answer should be:
a=9
b=12
c=15
test if this is valid:
9^2+12^2<=15^2
81+144<=225
255=255
Enjoy :)